### 25 Feb 2024 v1

@@ 4 Example use cases
Example use cases have been developed with OpenFOAM, in the release v2312555https://www.openfoam.com/download/release-history#v2312, and SmartSim release 0.6.2666https://github.com/CrayLabs/SmartSim/releases/tag/v0.6.2.
Implementation is publicly available on GitHub [^7] https://github.com/OFDataCommittee/openfoam-smartsim, as well as an archive of the project used to generate data in this manuscript [4].
@@ 4.1 Parameter tuning using ensembles of simulations
Many simulations have tuneable coefficients that affect their subgrid-scale parameterizations and numerics. Finding optimal sets of parameters that maximize a certain goodness of fit is directly analogous to hyperparameter optimization when training ML models. Like the ML problem, this problem is characterized by a large search space of $N$ dimensions (one for each tunable parameter) and having an objective function that is computationally costly (in ML, this is the cost of training the model; for simulation, this is the cost of running a model to completion).
Bayesian optimization (BO) attempts to mitigate both these issues by building a statistical model of the underlying objective function. The typical BO loop involves creating a prior distribution for the objective function, evaluating the function at select points, then using the results of the evaluation to construct a posterior distribution which can be used to generate the next set of query points. This has two primary advantages: 1) it does not need the derivatives of the actual function being evaluated, and 2) information from function evaluations improves the overall model. The former is particularly important for some numerical models for which an adjoint is not available. The latter is pertinent as well because shortening the time-to-solution is often difficult, but scaling the number of simulations performed in parallel is a simpler task
@@
Figure 9: The magnitude of the velocity from the large-eddy simulation (LES) of the Pitz and Daily case (top), the Reynolds-Averaged Navier Stokes using the default $k-\epsilon$ coefficients (middle), and the optimal coefficients (bottom). Note that for the LES simulation, the plotted field is the large-scale U that has had the eddy components filtered out.
@@ 4.2 Partitioned singular value decomposition
The SVD is a fundamental algorithm for analyzing and modeling fluid flows [8]. The flow’s state at time $t_{n}=n\Delta t$ may be expressed as vector $\mathbf{x}_{n}\in\mathbb{R}^{M}$. Based on the cell-centered finite volume method implemented in OpenFOAM, $\mathbf{x}_{n}$ may be composed of the cell-centered values of one or multiple fields, i.e., velocity, pressure, temperature, etc. A suitable normalization should be applied before building the state vector when mixing fields with different units [9]. The components of vector or tensor fields are concatenated. For example, the velocity field on mesh with $N_{p}$ cells yields a state vector of length $M=3N_{p}$. Given a time series of $N$ states, the state vectors are then arranged into a data matrix:
@@
While tremendously useful, memory consumption is a major challenge in the SVD’s application to fluid-mechanical problems. Consider a moderately sized mesh consisting of $N_{p}=10^{6}$ control volumes. Let us assume that $N=1000$ snapshots of velocity and pressure have been saved with double precision. Constructing the corresponding data matrix requires approximately 30GB of memory. Computing the economy SVD requires at least an additional 30GB. Hence, this computation could quickly outrun the memory of commodity workstations or compute nodes. A distributed memory parallel SVD implementation, e.g., the MPI-based variant available in ScaLAPACK, provides one workaround to handle the memory limitation. However, such implementations are not very well accessible in OpenFOAM or the typical ML ecosystem. Moreover, applying the SVD to state-of-the-art simulations in the automotive or aerospace sector can easily exceed the memory of an entire cluster.
@@
As before, the matrix multiplications (8) may be computed in parallel or sequentially. Moreover, it is not necessary to store the full SVD. Instead, we can analyze the singular values on the diagonal of $\mathbf{\Sigma}$, determine a suitable value for $r$, and keep the first $r$ columns of each matrix. Finally, we note that the rank-$r$ reconstruction to $\mathbf{X}$ can also be performed independently for each partition (sub-domain), i.e., $\mathbf{X}_{ri}=\mathbf{U}_{i}\mathbf{U}_{yi}\mathbf{\Sigma}_{y}\mathbf{V}^{T}_{y}$.
@@
    driver program: a Jupyter-Lab with Python kernel serves as the driver to start the SmartRedis database and to execute the steps listed below; to run OpenFOAM applications, we use SmartSim’s Experiment feature.
@@ 4.3.1 Problem definition
In this example use-case, we demonstrate how to leverage SmartSim to train and infer an ML model every time step in an OpenFOAM simulation on CFD data, effectively implementing the bidirectional CFD+ML workflow from fig. 2, for mesh motion, shown for a simplified example in fig. 11.
The mesh starts with an initial state fig. 10(a), deforms into an extreme state from fig. 10(b), then returns back to the initial state. This simplified use-case mirrors mesh motion, widely used in CFD in Fluid-Solid Interaction or Shape Optimization.
Mesh motion is implemented in OpenFOAM following OpenFOAM’s modular software design [3], making it possible to extend mesh motion with ML by developing a new mesh motion solver that uses an Artificial Neural Network (ANN) to approximate non-uniform mesh displacements, e.g., from fig. 11.
@@
For example, in OpenFOAM, there are mesh motion ‘engines’ (displacement maps), which define $M$ as an interpolation or a solution of a Laplace equation for $\mathbf{d}$ with variable diffusivity - the Laplacian mesh motion, given boundary displacements $\mathbf{d}|_{\partial\Omega}$.
@@ 4.3.2 Mesh motion in OpenFOAM
Since OpenFOAM’s parallel implementation relies on domain decomposition and MPI, boundary displacements are decomposed onto both boundary patches and over MPI ranks $r$. The decomposition of the domain boundary $\partial\Omega$ into boundary patches is necessary for applying different boundary conditions for (systems of) Partial Differential Equations (PDEs) solved in OpenFOAM. The decomposition of the boundary $\partial\Omega$ into patches such as the cylinder and outer wall in fig. 12, the domain decomposition for parallel program execution over $N_{r}$ MPI ranks, and time-marching in the CFD simulation over discrete time steps $\Delta t=t^{n+1}-t^{n}$, results in boundary displacements being split into sequences
@@
Note that it is quite possible that $\mathbf{d}^{n}_{p,r},\mathbf{p}^{n}_{p,r}=\emptyset$ (empty patches) if the domain decomposition results in the MPI rank $r$ not containing any part of the $p$-th boundary patch (e.g., the rank does not contain a part of the cylinder from fig. 12). This seemingly obvious case must be handled explicitly when exchanging boundary patch data with SmartRedis in the workflow (cf. fig. 2), as storing and reading zero-sized OpenFOAM fields in the database does not make sense, and creates unnecessary overhead. It is also important to note that mesh motion generally does not apply Neumann-type boundary conditions for the displacements at $\partial\Omega$; it uses Dirichlet-type (i.e., fixed-value) boundary conditions. The displacement map $M$ from eqs. 11 and 12 implemented in this CFD+ML algorithm as an ANN, views the Dirichlet (fixed value) boundary conditions simply as unstructured training data with boundary mesh points $\mathbf{p}|_{\partial\Omega}(t^{n})$ as features and the boundary mesh displacements $\mathbf{d}|_{\partial\Omega}(t^{n})$ as the labels of the ANN model of $M$. The decomposition given by eqs. 13 and 14 is not necessary, as long as discrete boundary displacements $\mathbf{d}^{n}|_{\partial\Omega}$ match the discrete boundary points $\mathbf{p}^{n}|_{\partial\Omega}$. Although the boundary decomposition in patches is not relevant for our mesh-motion example, we keep it to demonstrate how it impacts the communication and signaling workflow from fig. 2, in a general case where Neumann-type conditions might be necessary, e.g. for training a Physics-Informed Neural Network on OpenFOAM data with Neumann-type boundary conditions.
@@
The boundary decomposition from eqs. 13 and 14 increases the complexity of the CFD+ML workflow discussed in section 2.3 and shown in fig. 2. The CFD part of the OpenFOAM+SmartSim CFD+ML algorithm for mesh motion is summarized in pseudocode by algorithm 1. Boundary points and displacements that are used to train the $M$ as an ANN are aggregated over boundary patches for each MPI rank into SmartRedis datasets ($\mathbf{p}^{n+1}_{r}$ and $\mathbf{d}^{n+1}_{r}$), and the datasets are appended to respective points and displacements dataset lists, $L_{\mathbf{p}}$ and $L_{\mathbf{d}}$. This is done to simplify queries for received data in SmartRedis. Each individual MPI rank writes OpenFOAM data as SmartRedis tensors into SmartRedis using non-blocking communication for efficiency reasons. Another client of the SmartRedis database, in this example case, the ML training algorithm programmed in Python in SmartSim can easily query the size of the aggregation lists $L_{\mathbf{p,d}}$ until the target size $N_{r}$ number of MPI ranks - is reached and all data needed for the ANN training is available in SmartRedis. The OpenFOAM mesh motion solver, as a SmartRedis client, queries the SmartRedis database for the availability of $M$ as an ANN, trained on rank-agglomerated boundary points and displacements, $\mathbf{p}^{n+1}_{r}$ and $\mathbf{p}^{n+1}_{r}$, respectively. When $M$ becomes available, mesh displacements are evaluated at mesh points. There is a peculiarity in OpenFOAM at this point: mesh points in OpenFOAM are not stored in a geometric field [3] that would make it possible to differentiate between internal and boundary points easily. After forward inference is performed on all mesh points in SmartRedis, algorithm 1 must ensure that only internal displacements are overwritten. If we overwrote $\mathbf{d}^{n+1}|_{\partial\Omega}$, with approximated displacements, this would incur a loss of accuracy for exact displacements given by boundary conditions in OpenFOAM or more accurate displacements available as CFD boundary conditions in OpenFOAM. The overwriting of boundary conditions can be easily avoided in any other bidirectional CFD+ML algorithm, as is done here, by evaluating boundary conditions from OpenFOAM on fields inferred from an ML model.
@@ 4.3.3 Approximating mesh-motion displacements in SmartSim
We validate the example by comparing the non-orthogonality error [3] resulting from the mesh motion map $M$ approximated by the ANN in our CFD+ML algorithm, with the existing Laplacian mesh motion solver in OpenFOAM, which solves a Laplace equation for the displacements, in fig. 13. Non-orthogonality is a measure of mesh quality associated with face-centers; however, we visualize cell-averages of non-orthogonality angles. The hyperparameter tuning of the ANN used to model $M$ determines its accuracy, and in this example, we obtain similar distributions of the cell-average non-orthogonality angle to the one generated by the Laplacian mesh-motion solver. Still, the described approach to mesh-motion using the loosely-coupled bidirectional CFD+ML algorithms can easily be extended to improve the quality of the deformed mesh w.r.t. mesh motion solvers in OpenFOAM, e.g., by tuning the ANN hyperparameters or implementing derivative constraints to mesh motion, both of which we will address in our future work.
Algorithms 1 and 2 show the advantage of the loosely-coupled CFD+ML algorithms in SmartSim - besides the communication and signaling, each entity in a lossely coupled CFD+ML algorithm retains much of its original complexity.
@@ Acknowledgments
The authors gratefully acknowledge the organizing committee of the 18th OpenFOAM Workshop for the support provided during the preparation of this manuscript.
Andre Weiner gratefully acknowledges the Deutsche Forschungsgemeinschaft DFG (German Research Foundation) for funding parts of this work in the framework of the research unit FOR 2895 under the grant WE 6948/1-1.
Tomislav Maric gratefully acknowledges funding from July 1 2020 - 30 June 2024 by the German Research Foundation (DFG), Project-ID 265191195 - SFB 1194.
Alessandro Rigazzi and Andrew Shao were funded internally by the Hewlett Packard Enterprise HPC&AI business unit. They also thank the other developers of SmartSim, whose aid and support helped accelerate this work.

### 23 Apr 2024 v2

@@ 4 Example use cases
This section presents three separate use cases that emphasize the novel workflows that can be achieved using the SmartSim/OpenFOAM integration described above. The first case is an example of the AI-around-the-loop paradigm (where the AI algorithm interacts with the simulation on startup and finalization). Bayesian optimization is applied to ensembles of OpenFOAM simulations to identify specific values for tuneable parameters that maximize a desired metric. The second case implements a distributed, singular value decomposition algorithm as an example of online-analysis (analysis is performed on data streamed from the simulation in contrast to post-hoc analysis). The third case demonstrates an example of AI-in-the-loop (the AI technique is embedded as part of the solver). A neural network is first trained on boundary displacements and then immediately used to query displacements on the interior portion of the mesh
4.1 Data and software
Example use cases have been developed with OpenFOAM (version 2312 [2]), SmartRedis (version 0.5.2 [4]), and SmartSim (version 0.6.2 [3]). Both the software implementation and the starting files for the use case examples are publicly available on GitHub [5], as well as an archive of the project used to generate data in this manuscript [6].
@@ 4.2 Parameter tuning using ensembles of simulations
Many simulations have tunable coefficients that affect their subgrid-scale parameterizations and numerics. Finding optimal sets of parameters that maximize a certain goodness of fit is directly analogous to hyperparameter optimization when training ML models. Like the ML problem, this problem is characterized by a large search space of $N$ dimensions (one for each tunable parameter) and having an objective function that is computationally costly (in ML, this is the cost of training the model; for simulation, this is the cost of running a model to completion).
Bayesian optimization (BO) attempts to mitigate both these issues by building a statistical model of the underlying objective function [14]. The typical BO loop involves creating a prior distribution for the objective function, evaluating the function at select points, then using the results of the evaluation to construct a posterior distribution which can be used to generate the next set of query points. This has two primary advantages: 1) it does not need the derivatives of the actual function being evaluated, and 2) information from nearly any function evaluation improves the overall model. The former is particularly important for some numerical models for which an adjoint is not available. The latter is pertinent as well because shortening the time-to-solution is often difficult, but scaling the number of simulations performed in parallel is a simpler task.
@@
The $k-\epsilon$ turbulence closure traditionally has five free parameters. Based on physical arguments, this set can be reduced to three parameters [16], $C_{\mu}$, $C_{1}$, and $C_{2}$. The last free parameter is the dissipation rate $\epsilon$ at the inlet.
@@
Figure 9: The magnitude of the velocity from the large-eddy simulation (LES) of the Pitz and Daily case (a), the Reynolds-Averaged Navier Stokes using the default $k-\epsilon$ coefficients (b), and the optimal coefficients (c). Note that for the LES simulation, the plotted field is the large-scale U that has had the eddy components filtered out.
@@ 4.3 Partitioned singular value decomposition
SVD is a fundamental algorithm for analyzing and modeling fluid flows [19]. The flow’s state at time $t_{n}=n\Delta t$ may be expressed as vector $\mathbf{x}_{n}\in\mathbb{R}^{M}$. Based on the cell-centered finite volume method implemented in OpenFOAM, $\mathbf{x}_{n}$ may be composed of the cell-centered values of one or multiple fields, i.e., velocity, pressure, temperature, etc. A suitable normalization should be applied before building the state vector when mixing fields with different units [20]. The components of vector or tensor fields are then concatenated. For example, the velocity field on mesh with $N_{p}$ cells yields a state vector of length $M=3N_{p}$. Given a time series of $N$ states, the state vectors are then arranged into a data matrix:
@@
While tremendously useful, memory consumption is a major challenge in the application of SVD to fluid-mechanical problems. Consider a case with moderately sized mesh consisting of $N_{p}=10^{6}$ control volumes and $N=1000$ snapshots of velocity and pressure have been saved with double precision. Constructing the corresponding data matrix requires approximately 30GB of memory. Computing the economy SVD requires at least an additional 30GB. Even at relatively modest scales, his computation could quickly outrun the memory of commodity workstations or compute nodes. A distributed memory parallel SVD implementation, e.g., the MPI-based variant available in ScaLAPACK [21], provides one workaround to handle the memory limitation. However, such implementations are not very well accessible in OpenFOAM or the typical ML ecosystem. Moreover, applying the SVD to state-of-the-art simulations in the automotive or aerospace sector can easily exceed the memory of an entire cluster.
@@
As before, the matrix multiplications in equation (8) may be computed in parallel or sequentially. Moreover, it is not necessary to store the full SVD. Instead, we can analyze the singular values on the diagonal of $\mathbf{\Sigma}$, determine a suitable value for $r$, and keep the first $r$ columns of each matrix. Finally, we note that the rank-$r$ reconstruction to $\mathbf{X}$ can also be performed independently for each partition (sub-domain), i.e., $\mathbf{X}_{ri}=\mathbf{U}_{i}\mathbf{U}_{yi}\mathbf{\Sigma}_{y}\mathbf{V}^{T}_{y}$.
@@
    driver program: a Jupyter notebook with a Python kernel serves as the driver to start the SmartRedis database and to execute the steps listed below; to run OpenFOAM applications, we use SmartSim’s Experiment feature.
@@ 4.4.1 Problem definition
In this example use-case, we demonstrate how to leverage SmartSim to train an ML model and perform inference at every time step in an OpenFOAM simulation on CFD data, effectively implementing the bidirectional, AI-in-the-loop, CFD+ML workflow from fig. 2, for mesh motion, shown for a simplified example in fig. 11.
The mesh starts with an initial state fig. 10(a), deforms into an extreme state from fig. 10(b), then returns back to the initial state. This simplified use-case mirrors mesh motion, widely used in CFD applications involving fluid-solid interaction or shape optimization.
OpenFOAM implements mesh motion following its modular philosophy [12], making it possible to extend mesh motion with ML by developing a new mesh motion solver that uses an Artificial Neural Network (ANN) to approximate non-uniform mesh displacements, e.g., from fig. 11.
@@
For example, in OpenFOAM, there are mesh motion ‘engines’ (displacement maps), which define $M$ as an interpolation or a solution of a Laplace equation for $\mathbf{d}$ with variable diffusivity, i.e. the Laplacian mesh motion, given boundary displacements $\mathbf{d}|_{\partial\Omega}$.
@@ 4.4.2 Mesh motion in OpenFOAM
Since OpenFOAM’s parallel implementation relies on domain decomposition and MPI, boundary displacements are decomposed onto both boundary patches and over MPI ranks $r$. The decomposition of the domain boundary $\partial\Omega$ into boundary patches is necessary for applying different boundary conditions for systems of partial differential equations (PDEs) solved in OpenFOAM. The decomposition of the boundary $\partial\Omega$ into patches (i.e. the cylinder and outer wall in fig. 12), the domain decomposition over $N_{r}$ MPI ranks, and the time integration of the simulation over discrete time steps $\Delta t=t^{n+1}-t^{n}$, results in boundary displacements being split into sequences
@@
Note that it is quite possible that $\mathbf{d}^{n}_{p,r},\mathbf{p}^{n}_{p,r}=\emptyset$ (i.e. empty patches) if the domain decomposition results in the MPI rank $r$ not containing any part of the $p$-th boundary patch (e.g., the rank does not contain a part of the cylinder from fig. 12). This seemingly obvious case must be handled explicitly when exchanging boundary patch data with SmartRedis in the workflow (cf. fig. 2), as storing and reading zero-sized OpenFOAM fields in the database does not make sense, and creates unnecessary overhead. It is also important to note that mesh motion generally does not apply Neumann-type boundary conditions for the displacements at $\partial\Omega$; it uses Dirichlet-type (i.e., fixed-value) boundary conditions. The displacement map $M$ from eqs. 11 and 12 implemented in this CFD+ML algorithm as an ANN, views the Dirichlet (fixed value) boundary conditions simply as unstructured training data with boundary mesh points $\mathbf{p}|_{\partial\Omega}(t^{n})$ as features and the boundary mesh displacements $\mathbf{d}|_{\partial\Omega}(t^{n})$ as the labels of the ANN model of $M$. The decomposition given by eqs. 13 and 14 is not necessary, as long as discrete boundary displacements $\mathbf{d}^{n}|_{\partial\Omega}$ match the discrete boundary points $\mathbf{p}^{n}|_{\partial\Omega}$. Although the boundary decomposition in patches is not relevant for our mesh-motion example, we keep it to demonstrate how it impacts the communication and signaling workflow from fig. 2, in a general case where Neumann-type conditions might be necessary, e.g. for training a Physics-Informed Neural Network on OpenFOAM data with Neumann-type boundary conditions.
@@
The boundary decomposition from eqs. 13 and 14 increases the complexity of the CFD+ML workflow discussed in section 2.3 and shown in fig. 2. The CFD part of the OpenFOAM+SmartSim CFD+ML algorithm for mesh motion is summarized in pseudocode by algorithm 1. Boundary points and displacements that are used to train the $M$ as an ANN are aggregated over boundary patches for each MPI rank into SmartRedis datasets ($\mathbf{p}^{n+1}_{r}$ and $\mathbf{d}^{n+1}_{r}$), and the datasets are appended to respective points and displacements dataset lists, $L_{\mathbf{p}}$ and $L_{\mathbf{d}}$. This is done to simplify queries for available data in SmartRedis. Each individual MPI rank writes OpenFOAM data as SmartRedis tensors into SmartRedis using non-blocking communication for efficiency reasons. Another client of the SmartRedis database, in this example case, the python-based, ML training algorithm programmed can easily query the size of the aggregation lists $L_{\mathbf{p,d}}$ until the target size $N_{r}$ number of MPI ranks - is reached and all data needed for the ANN training is available in SmartRedis. The OpenFOAM mesh motion solver, as a SmartRedis client, queries the SmartRedis database for the availability of $M$ as an ANN, trained on rank-aggregated boundary points and displacements, $\mathbf{p}^{n+1}_{r}$ and $\mathbf{p}^{n+1}_{r}$, respectively. When $M$ becomes available, mesh displacements are evaluated at mesh points. There is a peculiarity in OpenFOAM at this point: mesh points in OpenFOAM are not stored in a geometric field [12] that would make it possible to differentiate between internal and boundary points easily. After forward inference is performed on all mesh points in SmartRedis, algorithm 1 must ensure that only internal displacements are overwritten. If we overwrote $\mathbf{d}^{n+1}|_{\partial\Omega}$, with approximated displacements, this would incur a loss of accuracy for exact displacements given by boundary conditions in OpenFOAM or more accurate displacements available as CFD boundary conditions in OpenFOAM. The overwriting of boundary conditions can be easily avoided in any other bidirectional CFD+ML algorithm, as is done here, by evaluating boundary conditions from OpenFOAM on fields inferred from an ML model.
@@ 4.4.3 Approximating mesh-motion displacements in SmartSim
We validate the example by comparing the non-orthogonality error [12] resulting from the mesh motion map $M$ approximated by the ANN in our CFD+ML algorithm, with the existing Laplacian mesh motion solver in OpenFOAM, which solves a Laplace equation for the displacements, in fig. 13. Non-orthogonality is a measure of mesh quality associated with face-centers; however, we visualize cell-averages of non-orthogonality angles. The hyperparameter tuning of the ANN used to model $M$ determines its accuracy. In this example, we have manually found a set of hyperparameters, which deliver similar distributions of the cell-average non-orthogonality angle to the ones generated by the Laplacian mesh-motion solver. In a real-world scenario, hyperparameter tuning would be integrated into the workflow, e.g., using the Bayesian Optimization from section 4.2. The described approach to mesh-motion using the loosely-coupled bidirectional CFD+ML algorithms can easily be extended to improve the quality of the deformed mesh w.r.t. mesh motion solvers in OpenFOAM, e.g., by tuning the ANN hyperparameters or implementing derivative constraints to mesh motion, both of which we will address in our future work.
Algorithms 1 and 2 show the advantage of the loosely-coupled CFD+ML algorithms in SmartSim - besides the communication and signaling, each entity in a loosely coupled CFD+ML algorithm retains much of its original complexity.
@@ Acknowledgments
The authors gratefully acknowledge the organizing committee of the 18th OpenFOAM Workshop for the support provided during the preparation of this manuscript. Andre Weiner gratefully acknowledges the Deutsche Forschungsgemeinschaft DFG (German Research Foundation) for funding parts of this work in the framework of the research unit FOR 2895 under the grant WE 6948/1-1.Tomislav Maric gratefully acknowledges funding from July 1 2020 - 30 June 2024 by the German Research Foundation (DFG), Project-ID 265191195 - SFB 1194.Alessandro Rigazzi and Andrew Shao were funded internally by the Hewlett Packard Enterprise HPC&AI business unit. They also thank the other developers of SmartSim, whose aid and support helped accelerate this work.